# Kolmogorov-size particles in homogeneous and isotropic turbulence

### v1

@@ 1.3 Present study
In this study, we investigate the fluid-solid interaction of suspension of Kolmogorov-size spherical particles moving in homogeneous isotropic turbulence at the relatively large microscale Reynolds number of $Re_{\lambda}\approx 140$ by use of direct numerical simulations. The study is based on both PR-DNS and PP-DNS. The objective of the present study is twofold; we aim (i) to investigate the modulation of forced homogeneous isotropic turbulence by finite Kolmogorov-size particles at a Reynolds number which is large enough to ensure a proper separation of scales, and (ii) to address the limits and the range of validity of the one-way-coupled PP-DNS in the simplest configuration of homogeneous isotropic turbulence. To do this, we consider a portion of the parameter space which is on the edge of the range of validity of the one-way-coupled PP-DNS (Brandt & Coletti, 2022).
The structure of the work is as follows. After this introduction, the computational set up and the numerical methods are described in §2. Then, section §3 is devoted to the assessment of the flow modulation, and discusses the results of the PR-DNS. The influence of the particles on the energy spectrum, on the scale-by-scale energy budget and on the local structure of the flow are discussed. Sections §4 and §5 deal respectively with the dynamics of the particles and with the inhomogeneity of their distribution in the flow. In these sections, we assess the validity of the one-way coupled PP-DNS. Eventually, concluding remarks are provided in §6.
@@ 2 Mathematical formulations and numerical method
where $\bm{u}=(u,v,w)$ is the fluid velocity, $p$ is the reduced pressure, and $\rho_{f}$ and $\nu$ are the fluid density and kinematic viscosity. At the right-hand-side of the momentum equation $\bm{f}$ is the Arnold-Beltrami-Childress (ABC) cellular forcing (Podvigina & Pouquet, 1994) used to inject energy at the largest scales and sustain turbulence, while $\bm{f}^{\leftarrow p}$ is the force the particles exert on the fluid phase.
@@ 2.1 Particle-resolved simulations (PR-DNS)
$\bm{f}^{\leftarrow f}=\oint_{\partial V_{p}}\bm{\sigma}\cdot\bm{n}\text{d}A,\ \bm{L}_{p}^{\leftarrow f}=\oint_{\partial V_{p}}\bm{r}\times\left(\bm{\sigma}\cdot\bm{n}\right)\text{d}A,$ (3)
@@ 2.2 One-way-coupled point particle simulations (PP-DNS)
The interface-resolved simulations are complemented with one-way-coupled point particle simulations (PP-DNS). In this work, we consider the complete governing equation for a point particle as introduced by Maxey & Riley (1983) and Gatignol (1983), i.e.
@@ 2.3 Computational details
We consider a single-phase micro-scale Reynolds number of $Re_{\lambda}=u^{\prime}\lambda/\nu\approx 140$ to ensure a relatively large inertial range of scales; here $u^{\prime}$ is the root mean square of the velocity fluctuations and $\lambda$ is the Taylor length scale. The particle diameter is set to $D_{p}/\eta\approx 0.9$, where $\eta$ is the Kolmogorov length scale for the single-phase case. Two volume fractions are considered, i.e. $\Phi_{V}=V_{s}/(V_{s}+V_{f})=10^{-5}$ and $\Phi_{V}=10^{-3}$ for a total number of particles of $N=742$ and $N=74208$, respectively. For each volume fraction two values of the particle density are considered, $\rho_{p}/\rho_{f}=5$ and $\rho_{p}/\rho_{f}=100$ to consider both light and heavy particles. This leads to a total of four PR-DNS. In PP-DNS the volume fraction is not a parameter, and only two simulations have been carried out for the different density ratios.
@@
For the PR-DNS the governing equations for the particles are dealt with the immersed boundary method introduced by Hori et al. (2022). The fluid-solid coupling is achieved in an Eulerian framework, and accounts for the inertia of the fictitious fluid inside the solid phase, so as to properly reproduce the particles’ behaviour in both the neutrally-buoyant case and in the presence of density difference between the fluid and solid phases. The soft sphere collision model (Cundall & Strack, 1979; Tsuji et al., 1993) is used to prevent the interpenetration between particles. A fixed-radius near neighbours algorithm (see Monti et al., 2021, and references therein) is used for the particle interaction to avoid an otherwise prohibitive increase of the computational cost when the number of particles becomes large. For the one-way-coupled PP-DNS the governing equation for the particle velocity is advanced in time using the second-order Adams-Bashforth time scheme. At each time step, the fluid velocity is evaluated at the position of the particle using a second-order linear interpolation.
@@
Details of the PR-DNS are reported in table 1. Note that, when looking at the bulk quantities, the flow modulation due to the solid phase is rather low.
[…]
$\Phi_{V}$ $\rho_{p}/\rho_{f}$ $N$ $\overline{\!\left\langle{\eta}\right\rangle\!}$ $\overline{\!\left\langle{E}\right\rangle\!}$ $\overline{\!\left\langle{\epsilon}\right\rangle\!}$ $St$ $Re_{p}$
@@ 3.1 Energy Spectrum
Figure 2: Energy spectrum for (top) $\Phi_{V}=10^{-5}$ and (bottom) $\Phi_{V}=10^{-3}$. The black line refers to the single-phase case, the red/blue light line is for $\rho_{p}/\rho_{f}=5$, an the red/blue dark line is for $\rho_{p}/\rho_{f}=100$. The symbols in the bottom panel are from the simulation carried out with the coarser grid. The filled circle on the $\kappa$ axis denotes the wavenumber associated with the particle size. The solid like is for the Kolmogorov $\kappa^{-5/3}$ scaling. The dashed line is for $\kappa^{-4}$.
@@
Figure 2 shows the influence of the solid phase on the energy spectrum $\mathcal{E}(\kappa)$, and highlights how Kolmogorov-size particles modulate the turbulent fluctuations scale-by-scale. The top and bottom panels are for $\Phi_{V}=10^{-5}$ and $\Phi_{V}=10^{-3}$, respectively. The black solid line refers to the reference unladen case; note the inertial range of scales where $\mathcal{E}\sim\kappa^{-5/3}$ extends for more than one decade of wavenumbers, confirming that the present Reynolds number is large enough to ensure a proper separation of scales. For validation purposes, in the bottom panel we also plot with symbols the energy spectrum obtained for the $\Phi_{V}=10^{-3}$ and $\rho_{p}/\rho_{f}=100$ case with the coarser grid, showing good agreement with those obtained with the standard grid, thus ensuring the suitability of the chosen grid resolution (see §2).
@@
At large scales, the spectra of the particle-laden cases substantially overlap with the unladen spectrum. We indeed observe only a weak depletion of the energy content at the intermediate scales; see the insets in the two panels. At scales smaller than a certain wavenumber $\kappa_{p}$, the energy spectra of the particle-laden cases deviate above the reference spectrum. Solid particles enhance the energy content of the small scales by means of their wake. Notably, this mechanism is amplified as $\Phi_{V}$ and/or $\rho_{p}/\rho_{f}$ increase, as conveniently visualised by the larger values of $\mathcal{E}$ for $\kappa>\kappa_{p}$ and by the shift of $\kappa_{p}$ towards smaller wavenumbers. However, notice that due to the low values of $\Phi_{V}$ considered, the flow modulation is rather low for all cases, being substantially negligible for $\Phi_{V}=10^{-5}$ and/or $\rho_{p}/\rho_{f}=5$.
@@
Figure 2 shows that the modulated energy spectrum exhibits multiscaling behaviour. The classical $\mathcal{E}(\kappa)\sim\kappa^{-5/3}$ decay in the inertial range of scales is indeed followed by a steeper decay $\mathcal{E}(\kappa)\sim\kappa^{-4}$ for $\kappa>\kappa_{p}$. A similar steep decay has been observed in bubbly flows at scales smaller than the bubble diameter, by means of both experiments (Martinez Mercado et al., 2010; Riboux et al., 2010; Prakash et al., 2016; Dung et al., 2023) and simulations (Pandey et al., 2020, 2022). Additionally, a similar multiscaling behaviour has also been observed in a turbulent planar Couette flow laden with small particles (Wang et al., 2023), and in homogeneous isotropic turbulence laden with slender fibres (Olivieri et al., 2022a, b). In the context of bubbly flows, the emergence of the $\kappa^{-\alpha}$ decay with $\alpha\geq 3$ has been attributed to the wakes the bubbles generate in a otherwise smooth flow (see for example Alméras et al., 2017; Pandey et al., 2020); here the velocity fluctuations produced by the bubbles are directly dissipated by viscosity (Lance & Bataille, 1991; Risso, 2018). Accordingly, this scaling has been indeed observed only when the bubble Reynolds number is large enough and is in the $10\leq Re_{bub}\leq 1000$ range (Pandey et al., 2020). To compare, we computed the local particle Reynolds number using the relative velocity between the particle and the surrounding flow, and found that it is in the $0.25\lessapprox Re_{p}\lessapprox 2.5$ range (see table 1). Further investigation on the link between the fluctuations induced by particles and the $\mathcal{E}(\kappa)\sim\kappa^{-4}$ decay is provided in §3.3 by looking at the scale-by-scale energy budget.
@@
Overall, figure 2 shows that particles almost do not modulate the inertial range of scales with the present parameters, where the classical energy cascade described by Richardson and Kolmogorov is preserved, but mainly affect the (otherwhise smooth) smallest scales of the flow.
@@ 3.2 Structure function and intermittency
Figure 3: Structure functions for (top) $\Phi_{V}=10^{-5}$ and (bottom) $\Phi_{V}=10^{-3}$. For each panel, from bottom to top the plots are for $S_{2}$, $S_{4}$ and $S_{6}$. The dashed lines represent $r^{p}$, while the dash-dotted ones $r^{p/3}$.
We extend the analysis done in the spectral domain by computing the longitudinal structure functions defined as $S_{p}(r)=\!\left\langle{\delta u(r)^{p}}\right\rangle\!$ where $\delta u(r)=(\bm{u}(\bm{x}+\bm{r})-\bm{u}(\bm{x}))\cdot\bm{r}/r$ and $r=|\bm{r}|$. In particular, figure 3 plots $S_{2}$, $S_{4}$ and $S_{6}$ as a function of $r$ for (top) $\Phi_{V}=10^{-5}$ and (bottom) $\Phi_{V}=10^{-3}$. In the single-phase case, we observe that $S_{p}\sim r^{p/3}$ in the inertial range of scales is in agreement with the Kolmogorov prediction (Kolmogorov, 1941), and that $S_{p}\sim r^{p}$ at the small scales, as a result of the differentiability of the fluid velocity field (Schumacher et al., 2007). Recall that although $S_{2}(r)$ is commonly referred to as scale energy (Davidson, 2004; Davidson & Pearson, 2005), its meaning slightly differs from $\mathcal{E}(\kappa)$; indeed while $\mathcal{E}(\kappa)\text{d}\kappa$ refers to the amount of energy associated with the scale $r=2\pi/\kappa$, $S_{2}(r)$ can be interpreted as the amount of energy associated with scales up to $r$. In agreement with the modulation of the energy spectrum, figure 3 shows that particles enhance the energy content at small scales compared to the unladen case. The energy enhancement is more intense for larger $\Phi_{V}$ and $\rho_{p}/\rho_{f}$, and becomes more evident when considering higher order structure functions. At the same time, the presence of the particles decreases the amount of energy stored at the largest scales (as seen by the blue and red curves laying below the black one). By interacting with the vortical structures of the flow, particles drain energy from scales larger than $D$ and reinject it back at smaller scales by means of their wake.

### v2

@@ 1.3 Present study
In this study, we investigate the fluid-solid interaction of suspension of Kolmogorov-size spherical particles moving in homogeneous isotropic turbulence at the relatively large microscale Reynolds number of $Re_{\lambda}\approx 140$ by use of direct numerical simulations. The study is based on both PR-DNS and one-way-coupled PP-DNS. The specific objective of the present study is twofold. We aim (i) to investigate (for the first time) the modulation of forced homogeneous isotropic turbulence by finite Kolmogorov-size particles at a Reynolds number which is large enough to ensure a proper separation of scales, and (ii) to address the limits and the range of validity of the one-way-coupled PP-DNS in the simplest configuration of homogeneous isotropic turbulence. To do this, we consider a portion of the parameter space which is on the edge of the range of validity of the one-way-coupled PP-DNS (Brandt & Coletti, 2022), and compare the results in terms of both Eulerian and Lagrangian particles’ statistics. Two volume fractions are considered, $\Phi_{V}=10^{-3}$ and $\Phi_{V}=10^{-5}$, and the solid-to-fluid density ratio is set equal to $\rho_{p}/\rho_{f}\approx 5$ and $\rho_{p}/\rho_{f}\approx 100$.
The structure of the work is as follows. After this introduction, the computational set up and the numerical methods are described in §2. Then, section §3 is devoted to the assessment of the flow modulation, and discusses the results of the PR-DNS. The influence of the particles on the energy spectrum, on the scale-by-scale energy budget and on the local structure of the flow are discussed. Sections §4 and §5 deal respectively with the dynamics of the particles and with the inhomogeneity of their distribution in the flow. In these sections, we assess the validity of the one-way-coupled PP-DNS. Eventually, concluding remarks are provided in §6.
@@ 2 Mathematical formulations and numerical method
where $\bm{u}=(u,v,w)$ is the fluid velocity, $p$ is the reduced pressure, and $\rho_{f}$ and $\nu$ are the fluid density and kinematic viscosity. At the right-hand-side of the momentum equation $\bm{f}$ is the Arnold-Beltrami-Childress (ABC) cellular forcing (Podvigina & Pouquet, 1994) used to inject energy at the largest scales and sustain turbulence, while $\bm{f}^{\leftarrow p}$ is the force the particles exert on the fluid phase. In this work $\bm{f}^{\leftarrow p}$ is non null for PR-DNS only.
@@ 2.1 Particle-resolved simulations (PR-DNS)
$\bm{f}^{\leftarrow f}=\oint_{\partial V_{p}}\bm{\sigma}\cdot\bm{n}\text{d}A\ \text{and}\ \bm{L}_{p}^{\leftarrow f}=\oint_{\partial V_{p}}\bm{r}\times\left(\bm{\sigma}\cdot\bm{n}\right)\text{d}A,$ (3)
@@ 2.2 One-way-coupled point particle simulations (PP-DNS)
The interface-resolved simulations are complemented with one-way-coupled point particle simulations (PP-DNS). Here the particles move under the action of the fluid flow, but they do not modulate it; the particle-particle interactions are also neglected. In this work, we consider the complete governing equation for a point particle as introduced by Maxey & Riley (1983) and Gatignol (1983), i.e.
@@ 2.3 Computational details
We consider a single-phase micro-scale Reynolds number of $Re_{\lambda}=u^{\prime}\lambda/\nu\approx 140$ to ensure a relatively large inertial range of scales; here $u^{\prime}$ is the root mean square of the velocity fluctuations and $\lambda$ is the Taylor length scale. The particle diameter is set to $D_{p}/\eta\approx 0.9$, where $\eta$ is the Kolmogorov length scale for the single-phase case. Two volume fractions are considered, i.e. $\Phi_{V}=V_{s}/(V_{s}+V_{f})=10^{-5}$ and $\Phi_{V}=10^{-3}$ for a total number of particles of $N=742$ and $N=74208$, respectively. For each volume fraction two values of the particle density are considered, $\rho_{p}/\rho_{f}=5$ and $\rho_{p}/\rho_{f}=100$ to consider both light and heavy particles. This leads to a total of four PR-DNS. In PP-DNS the particles do not modulate the flow and do not interact, therefore, the volume fraction is not a parameter. Because of this, only two PP-DNS simulations have been carried out for the different density ratios.
@@
For the PR-DNS the governing equations for the particles are dealt with by the immersed boundary method introduced by Hori et al. (2022). The fluid-solid coupling is achieved in an Eulerian framework, and accounts for the inertia of the fictitious fluid inside the solid phase, so as to properly reproduce the particles’ behaviour in both the neutrally-buoyant case and in the presence of density difference between the fluid and solid phases. The soft sphere collision model (Cundall & Strack, 1979; Tsuji et al., 1993) is used to prevent the interpenetration between particles and compute the $\bm{f}^{\leftrightarrow p}$ term in equation 2. In this model the particles are allowed to slightly penetrate. The collision is modelled as a spring and dashpot dynamical system, with the collision force being proportional to the penetration depth between the particles and to their relative normal velocity. A fixed-radius near neighbours algorithm (see Monti et al., 2021, and references therein) is used for the particle interaction to avoid an otherwise prohibitive increase of the computational cost when the number of particles becomes large. The PR-DNS simulations are performed without any additional lubrication correction, i.e. we consider only the lubrication that naturally arises from the method. For the one-way-coupled PP-DNS the governing equation for the particle velocity is advanced in time using the second-order Adams-Bashforth time scheme. At each time step, the fluid velocity is evaluated at the position of the particle using a second-order linear interpolation.
@@
Details of the PR-DNS are reported in table 1. Note that, when looking at the bulk quantities, the flow modulation due to the solid phase is rather low. For comparison, we mention that at a reference Reynolds number of $Re_{\lambda}\approx 240$ Hwang & Eaton (2006) experimentally report an energy attenuation of $\!\left\langle{E}\right\rangle\!/\!\left\langle{E_{0}}\right\rangle\!=0.78$ for a suspension of spherical particles with $D_{p}/\eta=0.9$, $\rho_{p}/\rho_{f}=2048$ and $\Phi_{V}\approx 10^{-4}$, which corresponds to a mass loading of $\Phi_{M}=\rho_{p}V_{p}/(\rho_{f}V_{f})=0.1$ (the $\cdot_{0}$ subscript refers to the single phase case). Although those parameters differ from ours, the value of the mass loading matches that of the $\rho_{p}/\rho_{f}=100$ and $\Phi_{V}=10^{-3}$ case ($\Phi_{M}\approx 0.098$), for which we however found a lower energy attenuation, i.e. $\!\left\langle{E}\right\rangle\!/\!\left\langle{E_{0}}\right\rangle\!=0.9\pm 0.26$. This suggests that the value of the mass loading alone is not sufficient to predict the level of the turbulence attenuation by Kolmogorov-size particles. For completeness, we mention that our predictions are close to those obtained by two-way-coupled PP-DNS at similar parameters, although at lower Reynolds numbers. For example, Boivin et al. (1998) predict an attenuation of $\!\left\langle{E}\right\rangle\!/\!\left\langle{E_{0}}\right\rangle\!\approx 0.88$ for $D_{p}/\eta=0.11$, $Re_{\lambda}=62$ and $\Phi_{M}=0.2$. At $Re_{\lambda}=38$, $D_{p}/\eta<1$, $\rho_{p}/\rho_{f}\gg 1$ and $\Phi_{M}=0.1$, instead, Squires & Eaton (1990) report an attenuation of $\!\left\langle{E}\right\rangle\!/\!\left\langle{E_{0}}\right\rangle\!=0.93$.
Table 1: Details of the PR-DNS considered in the present parametric study. $\epsilon$ is the dissipation rate. $St$ is the Stokes number defined as $St=\tau_{p}/\tau_{f}$, where $\tau_{p}=(\rho_{p}/\rho_{f})D_{p}^{2}/(18\nu)$ is the relaxation time of the particle and $\tau_{f}=\mathcal{L}/\sqrt{2\!\left\langle{E}\right\rangle\!/3}$ is the turnover time of the largest eddies. $Re_{p}$ is the particle Reynolds number defined as $Re_{p}=|\Delta\bm{u}|D/\nu$, where $\Delta\bm{u}=\bm{u}_{p}-\bm{u}_{f}$ is the fluid-particle relative velocity. Here, $\bm{u}_{f}$ is the fluid velocity seen by the particle evaluated as the average of the fluid velocity within a shell centred with the particle and with radius $R_{sh}=3(D_{p}/2)$ (see Uhlmann & Chouippe, 2017; Chiarini & Rosti, 2024).
$\Phi_{V}$ $\rho_{p}/\rho_{f}$ $N$ $\!\left\langle{\eta}\right\rangle\!$ $\!\left\langle{E}\right\rangle\!$ $\!\left\langle{\epsilon}\right\rangle\!$ $St$ $Re_{p}$
@@ 3.1 Energy Spectrum
Figure 2: Energy spectrum for (top) $\Phi_{V}=10^{-5}$ and (bottom) $\Phi_{V}=10^{-3}$. The black line refers to the single-phase case, the red/blue light line is for $\rho_{p}/\rho_{f}=5$, an the red/blue dark line is for $\rho_{p}/\rho_{f}=100$. The symbols in the bottom panel are from the simulation carried out with the coarser grid. The filled circle on the $\kappa$ axis denotes the wavenumber associated with the particle size. The solid line is for the Kolmogorov $\kappa^{-5/3}$ scaling. The dashed line is for $\kappa^{-4}$. Here $\kappa_{L}=2\pi/L$.
@@
Figure 2 shows the influence of the solid phase on the energy spectrum $\mathcal{E}(\kappa)$, and highlights how Kolmogorov-size particles modulate the turbulent fluctuations scale-by-scale. The top and bottom panels are for $\Phi_{V}=10^{-5}$ and $\Phi_{V}=10^{-3}$, respectively. The black solid line refers to the reference unladen case; note the inertial range of scales where $\mathcal{E}(\kappa)\sim\kappa^{-5/3}$ extends for more than one decade of wavenumbers, confirming that the present Reynolds number is large enough to ensure a proper separation of scales. For validation purposes, in the bottom panel we also plot with symbols the energy spectrum obtained for the $\Phi_{V}=10^{-3}$ and $\rho_{p}/\rho_{f}=100$ case with the coarser grid, showing good agreement with that obtained with the standard grid, thus ensuring the suitability of the chosen grid resolution (see §2). We compute the energy spectra (and the structure functions in the next section) using all the grid points of the computational domain including those that are inside the particles. However, as mentioned in Chiarini & Rosti (2024), with the present IBM the results do not change when the points within the particles are neglected.
@@
At large scales, the spectra of the particle-laden cases substantially overlap with the unladen spectrum. We indeed observe only a weak depletion of the energy content at the intermediate scales; see the insets in the two panels. At scales smaller than a certain wavenumber $\kappa_{p}$, the energy spectra of the particle-laden cases deviate above the reference spectrum. Solid particles enhance the energy content of the small scales by locally deforming the fluid flow around them. Notably, this mechanism is amplified as $\Phi_{V}$ and/or $\rho_{p}/\rho_{f}$ increase, as conveniently visualised by the larger values of $\mathcal{E}(\kappa)$ for $\kappa>\kappa_{p}$ and by the shift of $\kappa_{p}$ towards smaller wavenumbers. However, notice that due to the low values of $\Phi_{V}$ considered, the flow modulation is rather low for all cases, being substantially negligible for $\Phi_{V}=10^{-5}$ and/or $\rho_{p}/\rho_{f}=5$.
@@
Figure 2 shows that the modulated energy spectrum exhibits multiscaling behaviour. The classical $\mathcal{E}(\kappa)\sim\kappa^{-5/3}$ decay in the inertial range of scales is indeed followed by a steeper decay $\mathcal{E}(\kappa)\sim\kappa^{-4}$ starting at wave numbers close to $\kappa_{p}$. A similar steep decay has been observed in bubbly flows at scales smaller than the bubble diameter, by means of both experiments (Martinez Mercado et al., 2010; Riboux et al., 2010; Prakash et al., 2016; Dung et al., 2023) and simulations (Pandey et al., 2020, 2022). Additionally, a similar multiscaling behaviour has also been observed in a turbulent planar Couette flow laden with small particles (Wang et al., 2023), and in homogeneous isotropic turbulence laden with slender fibres (Olivieri et al., 2022a, b). In the context of bubbly flows, the emergence of the $\kappa^{-\alpha}$ decay with $\alpha\geq 3$ has been attributed to the wakes the bubbles generate in an otherwise smooth flow (see for example Alméras et al., 2017; Pandey et al., 2020); here the velocity fluctuations produced by the bubbles are directly dissipated by viscosity (Lance & Bataille, 1991; Risso, 2018). Accordingly, this scaling has been indeed observed only when the bubble Reynolds number is large enough and is in the $10\leq Re_{bub}\leq 1000$ range (Pandey et al., 2020). To compare, we computed the local particle Reynolds number using the relative velocity between the particle and the surrounding flow, and found that it is in the $0.25\lessapprox Re_{p}\lessapprox 2.5$ range (see table 1). More recently, Zamansky et al. (2024) observed that in bubbly flows the $\kappa^{-3}$ range constitutes a transition between the production-dominated (anisotropic) and the inertial/dissipative (isotropic) ranges of scales, and proposed that the specific $\alpha=3$ scaling results from the mean shear rate imposed by the bubbles which controls the rate of return to isotropy. In view of this, we mention that unlike what is commonly found in bubbly-induced agitation where the $\kappa^{-3}$ subrange is between the energetic and the inertial range of scales (see also figure 5 of Risso, 2018), in our case the $\kappa^{-4}$ scaling is observed at the small dissipative scales. It is also worth mentioning that in a recent work Ramirez et al. (2024) investigated the effect of singularities (e.g. induced by the particles) of various orders on the energy spectrum. They showed that singularities may affect the spectrum beyond the imposition of simple oscillations (Lucci et al., 2010), and actually cause power law scalings at wave numbers $\kappa\geq\kappa_{p}$, with the exact slope being dependent on the order of the singularity. However, in most of the cases investigated by us, the start of the $\kappa^{-4}$ scaling region appears at wave numbers sensibly lower than $\kappa_{p}$. Further investigation on the link between the fluctuations induced by particles and the $\mathcal{E}(\kappa)\sim\kappa^{-4}$ decay is provided in §3.3 by looking at the scale-by-scale energy budget.
@@
Overall, figure 2 shows that at the present parameters particles almost do not modulate the inertial range of scales, where the classical energy cascade described by Richardson and Kolmogorov is preserved, but mainly affect the (otherwise smooth) smallest scales of the flow.
@@ 3.2 Structure function and intermittency
Figure 3: Structure functions for (top) $\Phi_{V}=10^{-5}$ and (bottom) $\Phi_{V}=10^{-3}$. For each panel, from bottom to top the plots are for $S_{2}$, $S_{4}$ and $S_{6}$. The dashed lines represent $r^{p}$, while the dash-dotted ones $r^{p/3}$. The insets show a zoom of $S_{2}$ for $10\leq r/\eta\leq 40$.
We extend the analysis done in the spectral domain by computing the longitudinal structure functions defined as $S_{p}(r)=\!\left\langle{\delta u(r)^{p}}\right\rangle\!$ where $\delta u(r)=(\bm{u}(\bm{x}+\bm{r})-\bm{u}(\bm{x}))\cdot\bm{r}/r$ and $r=|\bm{r}|$. In particular, figure 3 plots $S_{2}$, $S_{4}$ and $S_{6}$ as a function of $r$ for (top) $\Phi_{V}=10^{-5}$ and (bottom) $\Phi_{V}=10^{-3}$. In the single-phase case, we observe that $S_{p}\sim r^{p/3}$ in the inertial range of scales is in agreement with the Kolmogorov prediction (Kolmogorov, 1941), and that $S_{p}\sim r^{p}$ at the small scales, as a result of the differentiability of the fluid velocity field (Schumacher et al., 2007). Recall that although $S_{2}(r)$ is commonly referred to as scale energy (Davidson, 2004; Davidson & Pearson, 2005), its meaning slightly differs from $\mathcal{E}(\kappa)$. Indeed while $\mathcal{E}(\kappa)\text{d}\kappa$ refers to the amount of energy associated with the scale $r=2\pi/\kappa$, $S_{2}(r)$ can be interpreted as the amount of energy associated with scales up to $r$. In agreement with the modulation of the energy spectrum, figure 3 shows that particles enhance the energy content at small scales compared to the unladen case. The energy enhancement is more intense for larger $\Phi_{V}$ and $\rho_{p}/\rho_{f}$, and becomes more evident when considering higher order structure functions. At the same time, for $\rho_{p}/\rho_{f}=100$ the presence of the particles decreases the amount of energy stored at the larger scales (as seen by the dark blue and dark red curves laying below the black one in the insets of figure 3). By interacting with the vortical structures of the flow, the (heavy) particles drain energy from scales larger than $D$ and reinject it back at smaller scales.